Having trouble understanding how to factor 4w^-25=0
I know how to factor out if there are three sets of numbers but with only two I cannot seem to get 4w^-25 to factor.

Answer by texttutoring(324)

I assume w^ means w^2.
4w^2-25=0
4w^2=25
w^2=25/4
Now take the square root of each side.
w =+/- (5/2)
which means w=5/2 or w=-5/2